Solve using the substitution method
3x+5y=19
X=33-7y
Having a hard time grasping the substitution method. Thanks!

Answer by Alan3354(69443)

Solve using the substitution method
3x+5y=19
x=33-7y
--------
The 2nd eqn is solved for x
Sub for x in the 1st eqn
-----
3x+5y=19
3(33-7y) + 5y = 19
Solve for y